October 24th
------------

The retroreflecting has remained off of the telescope because weather conditions were looking good (clear sky) at about 2:00PM


Night time activity
---------------------

The observations start at 18:30. there are clouds coming from west and seeing is around 2.5 arcsecond.The loop is closed using AOS succesfully around 19:30. However clouds do not allows to do reliable measurement of system performance because the CCD flux fluctuates too much. In some cases the star is lost because of the clouds. Due to the bad weather conditions we repeat some test with AOS that are succesfull. For about one hour we closed the loop on the Jupiter satellite Io that was found on a clear portion of the sky. The seeing values obscillates between 2 and 4 arcseconds. Around 20:30 we try to observe Jupiter using Io as reference star. This is a first test of using an off-axis object as reference star. The test was succesfull. However the seeing conditions were bad with values between 2 and 4 arcsec. We acquired some images of Jupiter with the AO loop closed but with intensity fluctuation on the CCD due to clouds and in bad seeing conditions. Around 23:30 the telescope is closed because of high umidity. The telescope remained close all night.

Test with dome closed.
----------------------

A test of the noise level in the WFS has been done with dome closed using the WFS internal reference source. The data should show what is the noise level in the WFS measurements vs. the number of incoming photons. Data reduction is ongoing.

SW test
-------

NOTE: The problem with TCS/IIF commands mentioned in yesterday report is relevant only for TCS build AO5.
Topic revision: r1 - 29 Oct 2010, AlfioPuglisi
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback